# Catching Stray Balls: Football, fandom, and the impact on digital discourse

### v1

@@ 5.1 Match Results and Shifts in Sentiment
Figure 2: Change in poster sentiment over 48 hour period around kick-off (FC Corpus).
@@ 5.2 Match Results and Post Habits
To analyse the relationship between match outcome and fan engagement, we subset data into posts made during a match (120-minutes) and up to 8-hours after, and calculated metrics to identify mean posts-per-match, the ratio between actual posts and expected posts, [^5] Post ratio is the percentage of total posts divided by the percentage of games for each outcome category. and average sentiment (Table 4 and 5). The analysis reveals how match outcomes influence both posting volume and sentiment over both time periods. [^6] Statistical significance was assessed in multiple ways. For sentiment analysis, one-way ANOVA tested overall differences in sentiment scores across match outcomes, followed by Tukey’s HSD post-hoc tests to identify specific group differences. For post counts $\chi$² goodness-of-fit tests were used to determine if frequencies differed significantly from those expected. The latter results were confirmed (p ¡ 0.001) with negative binomial regression models while controlling for season and club (subreddit) effects to account for temporal or club-specific variations that might skew aggregated results (especially as larger clubs, with more historic success, crowd out smaller clubs in the data).
@@
Table 4: Posts from all FC subreddits within 120-minutes of kick-off. $\chi$²-test for post distribution; ANOVA for sentiment differences.
@@
Table 5: Posts from all FC subreddits within 8-hours of kick-off. $\chi$²-test for post distribution; ANOVA for sentiment.
@@ 6 Cross-Community Sentiment Spread
Using a dataset of paired posts from club and non-club subreddits (the latter being posts made within 10-minutes of the former) we investigated how sentiment categories (negative, neutral, and positive) across post pairs are correlated (Kendall’s $\tau$). [^8] Kendall’s $\tau$ was used as it does not require data to be normally distributed. In our dataset, sentiment scores exhibit a trimodal distribution, with distinct peaks around negative, neutral, and positive values (see Table 1), which violates the normality assumption required by parametric tests (such as Pearson’s correlation coefficient). Additionally, as a non-parametric rank correlation measure, Kendall’s $\tau$ relies only on the ordering of the values, not their actual magnitudes (it evaluates whether sentiment scores in one context tend to align with those in another regardless of exact numerical differences, making it less susceptible to outliers or noisy data). In this context it assesses how consistently the relative ordering of one variable (e.g., sentiment in football posts) aligns with the ordering of another variable (e.g., sentiment in non-football posts). Additionally, it handles tied values better than Spearman’s $\rho$, which is important given that our sentiment data falls into three discrete categories (negative, neutral, positive) and has numerous tied rankings Kendall (1938). The analysis reveals a statistically significant — but very weak — correlation between post sentiment by the same users in FC and non-FC subreddits (Table 6). While a finding, this in itself is not surprising. One may expect a person’s emotional state to be consistent when posting across communities at similar times. However, our analysis shows the correlation doubles during matches (from $\tau$ = 0.059 to $\tau$ = 0.118). This suggests some transfer of football related emotional states to unrelated online spaces. That is, real-time football events influence sentiment in non-FC communities.
Following existing research showing that emotionally charged content spreads online more Brady et al. (2017), we also looked at these relationships with neutral posts made in FC subreddits removed (Table 7). Here correlations during matches strengthen further ($\tau$ = 0.118 to $\tau$ = 0.146), indicating emotionally charged users are more consistent in their emotional valence across community contexts, suggesting that heightened emotional states during matches may create stronger patterns of sentiment consistency in broader online behaviour.
@@ 8 Discussion and Conclusion
While this analysis focused on subreddits dedicated to specific football clubs — that is, forums for supporters of a specific club rather than general football related discussions — it should be noted that some of these football-related communities were retained in the dataset in Section 6. These are r/soccer (dedicated to general football discussion) and r/FantasyPL (dedicated to discussion around fantasy premier league football, such as a poster’s team, strategies, and results). While the statistical relationship remains with these subreddits removed (albeit with a diminished $\tau$), we reported results with them included for the following reasons.
@@ Appendix A Appendix: Subreddits
FC Subreddit Non-FC Subreddit Posts
@@ Appendix B Appendix: Post Pairs
Quotes and subreddits have been modified to avoid identification of authors.

### v2

@@ 5.1 Match Results and Shifts in Sentiment
Figure 2: Change in poster sentiment over 48-hour period around kick-off (FC Corpus).
@@ 5.2 Match Results and Post Habits
To analyse the relationship between match outcome and fan engagement, we subset data into posts made during a match (120-minutes) and up to 8-hours after, and calculated metrics to identify mean posts-per-match, the ratio between actual posts and expected posts, [^5] Post ratio represents relative posting intensity. It is the percentage of total posts for each outcome divided by the percentage of games for that outcome. This indicates whether posting activity is higher (¿1.0) or lower (¡1.0) than expected based on game frequency alone. and average sentiment (Table 4 and 5). The analysis reveals how match outcomes influence both posting volume and sentiment over both time periods. [^6] Statistical significance was assessed in multiple ways. For sentiment analysis, one-way ANOVA tested overall differences in sentiment scores across match outcomes, followed by Tukey’s HSD post-hoc tests to identify specific group differences. For post counts $\chi$² goodness-of-fit tests were used to determine if frequencies differed significantly from those expected. The latter results were confirmed (p ¡ 0.001) with negative binomial regression models while controlling for season and club (subreddit) effects to account for temporal or club-specific variations that might skew aggregated results (especially as larger clubs, with more historic success, crowd out smaller clubs in the data).
@@
Table 4: Posts from all FC subreddits within 120 minutes of kick-off. $\chi$²-test for post distribution; ANOVA for sentiment differences.
@@
Table 5: Posts from all FC subreddits within 8 hours of kick-off. $\chi$²-test for post distribution; ANOVA for sentiment.
@@ 6 Cross-Community Sentiment Spread
Using a dataset of paired posts from club and non-club subreddits (the latter being posts made within 10-minutes of the former) we investigated how sentiment categories (negative, neutral, and positive) across post pairs are correlated (Kendall’s $\tau$). [^8] Kendall’s $\tau$ was used as it does not require data to be normally distributed. In our dataset, sentiment scores exhibit a trimodal distribution, with distinct peaks around negative, neutral, and positive values (see Figure 1), which violates the normality assumption required by parametric tests (such as Pearson’s correlation coefficient). Additionally, as a non-parametric rank correlation measure, Kendall’s $\tau$ relies only on the ordering of the values, not their actual magnitudes (it evaluates whether sentiment scores in one context tend to align with those in another regardless of exact numerical differences, making it less susceptible to outliers or noisy data). In this context it assesses how consistently the relative ordering of one variable (e.g., sentiment in football posts) aligns with the ordering of another variable (e.g., sentiment in non-football posts). Additionally, it handles tied values better than Spearman’s $\rho$, which is important given that our sentiment data falls into three discrete categories (negative, neutral, positive) and has numerous tied rankings Kendall (1938). The analysis reveals a statistically significant — but very weak — correlation between post sentiment by the same users in FC and non-FC subreddits (Table 6). While a finding, this in itself is not surprising. One may expect a person’s emotional state to be consistent when posting across communities at similar times. However, our analysis shows the correlation doubles during matches (from $\tau$ = 0.059 to $\tau$ = 0.118). This suggests some transfer of football related emotional states to unrelated online spaces. That is, real-time football events influence sentiment in non-FC communities.
Following existing research showing that emotionally charged content spreads online more quickly Brady et al. (2017), we also looked at these relationships with neutral posts made in FC subreddits removed (Table 7). Here correlations during matches strengthen further ($\tau$ = 0.118 to $\tau$ = 0.146), indicating emotionally charged users are more consistent in their emotional valence across community contexts, suggesting that heightened emotional states during matches may create stronger patterns of sentiment consistency in broader online behaviour.
@@ 8 Discussion and Conclusion
While this analysis focused on subreddits dedicated to specific football clubs — that is, forums for supporters of a specific club rather than general football related discussions — it should be noted that some of these football-related communities were retained in the dataset used in Section 6. These are r/soccer (dedicated to general football discussion) and r/FantasyPL (dedicated to discussion around fantasy premier league football, such as a poster’s team, strategies, and results). While the statistical relationship remains with these subreddits removed (albeit with a diminished $\tau$), we reported results with them included for the following reasons.
@@ Appendix A Appendix: Subreddits
Subreddit Football Club Posts